A line contains the points (8,y) and (11,7) and has a slope of 6.find the value of y.
Answer by mananth(16946)

slope of a line formula when co ordinates of two points are given
(y2-y1)/(x2-x1)=m
(7-y)/(11-8) =6
(7-y)= 6*3
7-y=18
y=-11